# Exocentric To Egocentric Transfer For Action Recognition: A Short Survey

### v1

@@ III-F Exo-Ego Transfer
A vast amount of knowledge in the form of motion cues is embedded in exocentric videos that can be transferred to the egocentric domain. Ardeshir et al. [26] is a premier work that learns mappings between the ego-exo views. In Ardeshir et al. [96], the authors propose a two-stream view-specific architecture to adapt from exo to ego view. Ho et al. [97] utilizes a semi-supervised domain adaptation technique to adapt exocentric visual cues to egocentric videos. Xu et al. [98] uses a prompt-masking technique for transferring information for egocentric hand-object interaction. Different from previous approaches, Li et al. [25] proposes an improved pre-training approach to extract signals from exocentric videos helpful for the egocentric domain. Ohkawa et al. [99] aids further adaptation by performing view-invariant pretraining and finetuning. Different from previous techniques, Quattrocchi et al. [100] proposes an adaptation technique for temporal action segmentation.
@@
In Nishimura et al. [101], the geometric transformation is used to tackle a novel problem of view-verification (bird’s eye-view trajectory estimation) is computed from the egocentric movement. Qian et al. [102] is an extension to a more challenging problem of bird’s eye view estimation in the absence of proper calibration.
III-G Joint ego-exo works
@@
This section outlines works that aim to learn a joint ego-exo representation. Sigurdsson et al. [103] makes the first attempt to jointly relate first-person and third-person viewpoints. Yu et al. [104, 105] leverages a joint attention mechanism to extract a shared representation between the views. In Wang et al. [106], a sentence-bert language model [107] is utilized to semantically align the unpaired exocentric and egocentric videos. Xue et al.[42] became the first work to propose a self-supervised learning approach to learn a view-invariant representation. Zhao et al. [108] solves a novel task of identifying and segmenting the egocentric camera wearer in a third-person view.
@@
3D egocentric pose estimation has also benefited from a joint ego-exo learning framework [109, 110]. A novel thermal image-based 3D hand-pose dataset has been released in ThermoHands [111]. Lu et al. [112] covers a scene-graph generation technique based on a self-attention mechanism between the ego and exo views. The authors of Wen et al. [113] present a solution combining third-person and first-person images to predict the subject’s location from the third-person viewpoint. Jia et al. [114] extracts exocentric and egocentric conversational signals to generate a scene graph. Xu et al. [115] shows an improvement in egocentric captioning by retrieving semantically relevant third-person videos [116].
III-H Miscellaneous applications
@@
IV Conclusion

### v2

@@ 3.6 Exo-Ego Transfer
A vast amount of knowledge in the form of motion cues is embedded in exocentric videos that can be transferred to the egocentric domain. Transferring knowledge from exocentric to egocentric views present several key challenges, primarily due to significant viewpoint mismatches, occlusions, and differences in motion representation. Ardeshir et al. [92] is a premier work that learns mappings between the ego-exo views. In Ardeshir et al. [93], the authors propose a two-stream view-specific architecture to adapt from exo to ego view. Ho et al. [94] utilizes a semi-supervised domain adaptation technique to adapt exocentric visual cues to egocentric videos. Xu et al. [95] uses a prompt-masking technique for transferring information for egocentric hand-object interaction. These domain adaptation techniques require expensive large scale-datasets. Unlike earlier methods, Li et al. [91] introduce an enhanced training strategy that effictively extracts signals from exocentric videos for application in the egocentric domain, as shown in Figure 9. Ohkawa et al. [96] aids further adaptation by performing view-invariant pretraining and finetuning. Different from previous techniques, Quattrocchi et al. [97] proposes an adaptation technique for temporal action segmentation.
@@
In Nishimura et al. [98], the geometric transformation is used to tackle a novel problem of view-verification (bird’s eye-view trajectory estimation) is computed from the egocentric movement. However, this approach may encounter challenges in real-world scenarios where the camera placement is inconsistent. Qian et al. [99] is an extension to a more challenging problem of bird’s eye view estimation in the absence of proper calibration.
3.7 Joint Ego-Exo Works
Figure 10: Temporally aligned view-invariant representation learned from the ego-exo videos, as proposed by [35].
@@ 3.7 Joint Ego-Exo Works
This section outlines works that aim to learn a joint ego-exo representation. Sigurdsson et al. [100] makes the first attempt to jointly relate first-person and third-person viewpoints. Yu et al. [101, 102] leverages a joint attention mechanism to extract a shared representation between the views. These approaches are reliant on learning an attention model, making them computationally intensive. In Wang et al. [103], a sentence-bert language model [104] is utilized to semantically align the unpaired exocentric and egocentric videos. Xue et al. [35] became the first work to propose a self-supervised learning approach to learn a view-invariant representation. Figure 10 illustrates the approach. Zhao et al. [105] solves a novel task of identifying and segmenting the egocentric camera wearer in a third-person view.
@@
3D egocentric pose estimation has also benefited from a joint ego-exo learning framework [106, 107]. A novel thermal image-based 3D hand-pose dataset has been released in ThermoHands [108]. Lu et al. [109] covers a scene-graph generation technique based on a self-attention mechanism between the ego and exo views. The authors of Wen et al. [110] present a solution combining third-person and first-person images to predict the subject’s location from the third-person viewpoint. Jia et al. [111] extracts exocentric and egocentric conversational signals to generate a scene graph. Xu et al. [112] shows an improvement in egocentric captioning by retrieving semantically relevant third-person videos [113]. Future research should explore more robust multi-view fusion techniques, better temporal modeling, and efficient, scalable approaches to address these challenges and unlock new possibilities for joint ego-exo applications.
3.8 Miscellaneous Applications
@@
In recent years, VLMs have shown their effectiveness on various vision tasks, such as action anticipation [133, 134], anomaly detection [135, 136], and affective understanding [137, 138]. Existing VLMs [139, 140] trained on large amounts of third-person perspective contain many signals that can be transferable for egocentric tasks. Previous works like [91, 92] utilize simpler architectures to learn egocentric representations from third-person data. VLMs would be more capable of learning stronger representations for more complicated joint ego-and-exo tasks.
4 Conclusion and Future Directions
@@
Given the evolving research landscape, we propose several promising directions for future exploration in joint egocentric and exocentric applications. The details are shown below.
  1. 1.
    Fine-Grained Multi-View Correspondences: A major challenge in joint ego-exo tasks is the accurate alignment across the first and third-person views. The viewpoint shift between ego and exo view introduces occlusions, geometric distortions, and ambiguities in the hand-object interactions. Current methods, reliant on feature-based similarity matching struggle when the viewpoints observe different object appearances and motion trajectories. Future research should explore 3D-aware representations to better model fine-grained correspondences between the egocentric and exocentric data.
  2. 2.
    Temporal dependencies across perspectives: Existing works primarily focus on spatial relationships, while the temporal consistency acoss ego-exo views remains unexplored. Recent studies [91] indicate the potential of temporal alignment techniques to enhance predictive modeling of long-term interactions and actions.Investigating self-supervised sequence alignment could enhance predictive long-term predictive modelling.
  3. 3.
    Robustness to noisy, misaligned, and adversarial inputs: Practical deployment of ego-exo models in applications such as sports analytics, autonomous systems, and assistive AI requires robustness to real-world variations. Current models fail when faced with sensor noise, occlusions, dynamic lighting, and adversarial perturbations. Developing adaptive fusion mechanisms, self-supervised robustness objectives, or adversarial training tailored for multi-view learning could significantly enhance model generalization. Additionally, leveraging multi-modal cues (e.g., audio, depth, IMU data) could provide redundancy and improve stability in challenging environments.
  4. 4.
    Realtime and Efficient Processing: Current models are computationally expensive, limiting real-world deployment in robotics, AR/VR and autonomous systems.
  5. 5.
    Dataset Collection and Annotation: There is a need for even richer datasets that capture the diverse multi-view scenarios across challenging environments. Additionally, capturing synchronized ego-exo data at a higher frame rate could facilitate better temporal alignment.
  6. 6.
    Multi-modal fusion and sensor integration: Incorporating multi-modal data such as sensors, IMUs, and LiDAR can enhance robustness to occlusions, varying lightning conditions, and motion blur.
By addressing these challenges, future research can enable more generalizable, temporally aware, pratical, and robust ego-exo models, unlocking new possibilities for human activity recognition, immersive AR/VR, and robotics.